# DREAMS: A python framework to train deep learning models with model card reporting for medical and health applications

### v1

@@ Abstract
Electroencephalography (EEG) data provides a non-invasive method for researchers and clinicians to observe brain activity in real time. The integration of deep learning techniques with EEG data has significantly improved the ability to identify meaningful patterns, leading to valuable insights for both clinical and research purposes. However, most of the frameworks so far, designed for EEG data analysis, are either too focused on pre-processing or in deep learning methods per, making their use for both clinician and developer communities problematic. Moreover, critical issues such as ethical considerations, biases, uncertainties, and the limitations inherent in AI models for EEG data analysis are frequently overlooked, posing challenges to the responsible implementation of these technologies. In this paper, we introduce a comprehensive deep learning framework tailored for EEG data processing, model training and report generation. While constructed in way to be adapted and developed further by AI developers, it enables to report, through model cards, the outcome and specific information of use for both developers and clinicians. In this way, we discuss how this framework can, in the future, provide clinical researchers and developers with the tools needed to create transparent and accountable AI models for EEG data analysis and diagnosis.
@@
1 Motivation and significance
Existing frameworks dedicated to analyzing EEG data are relatively scarce and those that do exist tend to either cover a wide range of topics inadequately or prioritize either the initial steps of data preparation or the utilization of advanced machine learning techniques [1]. Preprocessing-centric frameworks offer tools tailored to tasks such as noise removal, artifact correction and data segmentation, but may lack advanced analysis capabilities [2]. Deep learning-centric frameworks emphasize using sophisticated algorithms for tasks like classification and feature extraction, assuming preprocessed data [3]. Indeed, this lack of comprehensive frameworks presents challenges for researchers and clinicians seeking effective tools for EEG data analysis.
Furthermore, in the realm of EEG data analysis, ethical considerations play a critical role [4], requiring strict adherence to ethical standards throughout the process of data collection and analysis. This includes securing informed consent from participants and safeguarding their privacy. Biases [5] represent a crucial concern in EEG data analysis, as they can distort results if left unaddressed. These biases, which include selection bias and interpretation bias, must be carefully managed to ensure accurate findings. Uncertainties arising from the complexity of EEG signals present another significant factor that requires acknowledgment and quantification to ensure reliable interpretation. AI models utilized in EEG analysis have inherent limitations, such as interpretability issues, which must be accounted for to prevent misleading conclusions and ensure responsible research and clinical practices [6]. Unfortunately, it is all too common for these critical considerations to be disregarded or given insufficient attention.
Without proper documentation, it becomes difficult to evaluate reliability of the models, effectiveness and potential biases, which are critical factors in ensuring safe and ethical use.
To address the challenges above, we introduce Deep REport for AI ModelS (DREAMS), a Python framework for documenting the performance characteristics of trained machine learning (ML) and deep learning (DL) models. Table 1 provides the main metadata of the code of DREAMS.
@@ 2 Related work
Figure 1: Overview of DREAMS overall architecture.
@@ 3.2.2 Model training and evaluation
At this stage, we meticulously document all the paths to vital information related to both training and evaluation within the YAML file. This DREAMS comprehensive record encompasses not only paths to graphics but also to other essential data, ensuring that every critical component is easily accessible and well-organized. This approach ensures efficient management and retrieval of information throughout the training and evaluation processes.
@@ 3.2.3 Model card generation
An example of a snippet code using generate_modelcard() function is as follows:
@@
4 An illustrative example for EEG data analysis
To exemplify our approach, we utilize the training pipeline of a deep learning model in conjunction with the Finer-grained Affective Computing EEG Dataset (FACED) dataset [11], which is compatible with TorchEEG  [^1] https://torcheeg.readthedocs.io/en/latest/.
@@ 4 An illustrative example for EEG data analysis
The EEG data were collected from 123 subjects using 32 electrodes positioned according to the international 10–20 system. During the recording sessions, participants viewed 28 different video clips designed to elicit various emotional responses. These video clips spanned nine distinct emotion categories: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral emotion.
@@
The model described in this example classifies EEG data into three main categories: negative (0), which includes emotions such as anger, disgust, fear, and sadness; neutral (1); and positive (2), which includes emotions like amusement, inspiration, joy and tenderness.
To prepare the data for this classification, it is processed through online transformation into tensors and a 2222D format. Then, the labels are processed by adjusting their values accordingly. Following this preprocessing, the dataset is divided into training and validation sets. Subsequently, DataLoaders are created for each set, facilitating the management and batching of data for the training and validation processes. Implementing these steps can be as follows:
@@
3dataset = FACEDDataset(root_path= data_folder,
@@
Next, we set up the TSception model [12] and define its hyperparameters:
1#Initiate the model
2model = TSCeption(num_classes=9,
3 num_electrodes=30,
4 sampling_rate=250,
5 num_T=15,
6 num_S=15,
7 hid_channels=32,
8 dropout=0.5)
TSception is a multi-scale convolutional neural network designed to classify emotions from EEG data. It incorporates dynamic temporal layers, asymmetric spatial layers and high-level fusion layers, enabling the model to learn discriminative representations across both time and channel dimensions simultaneously.
Following this, we proceed to configure the training and validation processes utilizing PyTorch. This involves setting up the necessary parameters and components for training the model, such as defining the loss functions, optimizers, and evaluation metrics. We establish the procedures for validating the model’s performance, including managing the data flow, monitoring metrics, and adjusting the training process as needed to ensure effective learning and evaluation. A standard implement for the training and validation configuration is:
@@
Upon completion of these procedures, we utilize a custom plotting function to create and store plots depicting accuracy, loss and the confusion matrix in the specified directory:
@@
3plot_confidence_intervals(precision=results[’precision’], recall=results[’recall’],
4 accuracy=results[’accuracy’], f1=results[’f1_score’],
5 n=len(val_loader.dataset), save_path="./logs/CI_plot.png",
6 confidence=0.95)
Lastly, we call the generate_model_card() function, providing it with the path to the YAML configuration file  [^2] https://github.com/LucidJun/DREAMS/tree/main/template, the output path for saving the model card, and the version number of the model card:
@@
Figure 2 provides a detailed overview of the generated model card for this example. This model card encapsulates the key information and metrics resulting from the model’s training and evaluation processes, offering a comprehensive summary of the model’s performance, data characteristics and other relevant details.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Model cards generated for the example described in Sec. 4.
5 Impact

### v2

# DREAMS: A python framework for Training Deep Learning Models on EEG Data with Model Card Reporting for Medical Applications

@@ Abstract
Electroencephalography (EEG) provides a non-invasive way to observe brain activity in real time. Deep learning has enhanced EEG analysis, enabling meaningful pattern detection for clinical and research purposes. However, most existing frameworks for EEG data analysis are either focused on preprocessing techniques or deep learning model development, often overlooking the crucial need for structured documentation and model interpretability. In this paper, we introduce DREAMS (Deep REport for AI ModelS), a Python-based framework designed to generate automated model cards for deep learning models applied to EEG data. Unlike generic model reporting tools, DREAMS is specifically tailored for EEG-based deep learning applications, incorporating domain-specific metadata, preprocessing details, performance metrics, and uncertainty quantification. The framework seamlessly integrates with deep learning pipelines, providing structured YAML-based documentation. We evaluate DREAMS through two case studies: an EEG emotion classification task using the FACED dataset and a abnormal EEG classification task using the Temple Univeristy Hospital (TUH) Abnormal dataset. These evaluations demonstrate how the generated model card enhances transparency by documenting model performance, dataset biases, and interpretability limitations. Unlike existing model documentation approaches, DREAMS provides visualized performance metrics, dataset alignment details, and model uncertainty estimations, making it a valuable tool for researchers and clinicians working with EEG-based AI. The source code for DREAMS is open-source, facilitating broad adoption in healthcare AI, research, and ethical AI development.
@@
1 Introduction
Electroencephalography (EEG)-based deep learning has demonstrated significant potential in various applications, including brain-computer interfaces (BCI), emotion recognition, and neurological disorder diagnosis [1, 2, 3]. Despite these advancements, a key challenge in the development and deployment of deep learning models for EEG applications is the lack of structured and standardized model documentation. Models designed for EEG based applications are highly sensitive to variations in dataset preprocessing, feature extraction techniques, and model architectures, making comparability and reproducibility challenging [4].
Reproducibility is a fundamental requirement in healthcare applications of machine learning (ML) models, yet EEG-based deep learning research and applications continues to suffer from inconsistent and non-transparent model reporting practices. Many studies report only high-level performance metrics (e.g., accuracy, F1-score) while omitting crucial details such as preprocessing pipelines, hyperparameter configurations, and prediction uncertainties. The absence of structured documentation not only hinders reproducibility but also complicates model benchmarking and comparative analysis, limiting the ability of researchers to validate findings across different datasets and experimental settings. This challenge becomes even more critical in clinical applications, where a lack of reproducibility can impede regulatory approval, deployment, and real-world patient outcomes. Addressing these gaps requires a standardized and automated model documentation framework that ensures transparent, structured, and reproducible reporting of EEG-based deep learning models.
One of the key challenges in deploying deep learning models is their inherent uncertainty in predictions, often leading to overconfidence in incorrect classifications [5, 6]. This issue is particularly critical in EEG-based deep learning, where high signal variability, inter-subject differences, and noise hinder model generalization capacity [7]. In clinical settings, failing to quantify uncertainty can result in misdiagnoses or incorrect treatments, posing high risks [4].
Uncertainty estimation enhances risk-aware decision-making, allowing models to flag low-confidence predictions for further review instead of making potentially harmful automatic decisions. For example, an EEG-based seizure detection system with uncertainty awareness can alert clinicians when a prediction is inconclusive, prompting manual verification. Integrating uncertainty estimation into structured model documentation, such as model cards, ensures greater transparency, interpretability, and trustworthiness in EEG-based healthcare AI [8].
Figure 1: Overview of the potential role and usefulness of DREAMS for EEG data analysis in the context of clinical research.
To address these challenges, we introduce DREAMS, a Python-based model card framework specifically designed for EEG-based deep learning models. DREAMS provides automated documentation covering key aspects such as dataset characteristics, preprocessing techniques, model architecture, performance evaluation with confidence intervals, and version tracking.
Our main contributions are:
  1. 1.
    We propose DREAMS, a structured model card framework for EEG deep learning, enabling comprehensive documentation of model performance, preprocessing pipelines, and uncertainty estimation.
  2. 2.
    We introduce confidence interval-based uncertainty reporting, improving model interpretability and helping researchers assess the prediction reliability of the models.
  3. 3.
    We automate model documentation with version tracking, allowing for efficient benchmarking and reproducibility across different model versions and EEG datasets.
  4. 4.
    We demonstrate the framework’s applicability in EEG-based classification tasks, showcasing its potential for broader EEG-based ML applications.
@@ 1 Introduction
Figure 1 sketches DREAMS role in the context of clinical research based in EEG data. Table 1 provides the main metadata of the code of DREAMS.
@@ 2 Related work
2.1 Comparison with Existing Model Documentation Frameworks
The increasing adoption of deep learning in EEG analysis necessitates structured documentation tools that ensure transparency, reproducibility, and ethical compliance. While existing model documentation frameworks offer general-purpose solutions, they often lack domain-specific adaptability for EEG-based AI. This section compares DREAMS with prominent model documentation approaches, highlighting its unique contributions.
Several model documentation framework exists, but they are primarily designed for general machine learning models rather than the specialized needs of EEG-based deep learning. One such framework is the Google’s Model Card Toolkit (MCT) [12] that provides structured model reporting, but it is heavily integrated with TensorFlow Extended (TFX) and lacks direct support for EEG preprocessing metadata or domain-specific performance metrics. Similarly, Hugging Face Model Cards [13] have become a standard for documenting NLP and computer vision models, but they do not accommodate the interpretability challenges, uncertainty quantification, or ethical considerations unique to EEG-based deep learning. While the traditional manual model documentation methods [9] (e.g., static reports, research papers) is flexible but lacks automation, standardization, and reproducibility, making it time-consuming and inconsistent across studies. Without a structured and automated approach, documenting EEG-based deep learning models remains inefficient and prone to inconsistencies, underscoring the need for a dedicated framework like DREAMS.
Unlike these frameworks, DREAMS is specifically designed for EEG-based deep learning models and offers several key advantages, namely:
  • EEG-Specific Adaptation. DREAMS includes EEG preprocessing details, dataset alignment reports, and uncertainty estimation, ensuring domain-relevant transparency.
  • Automated Documentation. Unlike manual reporting, DREAMS automatically logs model parameters, training details, and performance metrics, reducing human error and improving reproducibility.
  • Framework Agnostic. Unlike Google’s Model Card Toolkit, which is tied to TensorFlow, DREAMS works with any deep learning framework (e.g., PyTorch, Keras, TensorFlow).
  • Enhanced Interpretability. DREAMS provides visualized model performance metrics, confidence intervals, and dataset biases, making it more informative than text-heavy model cards.
  • Ethical and Regulatory Compliance. By documenting bias risks, dataset composition, and interpretability challenges, DREAMS aligns with best practices in responsible AI for medical applications.
Table 2 summarizes the differences between DREAMS and existing model documentation approaches.
Feature Google MCT Hugging Face Manual Docs DREAMS
Designed for EEG Models? ✗ (No) ✗ (No) (Limited) ✓ (Yes)
Automated Model Card Generation (Partial, TFX) ✗ (No) ✗ (No) ✓ (Yes)
Integration with DL Frameworks TensorFlow Only PyTorch /TensorFlow ✗ (No) Works with Any DL Framework
Dataset Transparency Basic Metadata Basic Metadata (Manual) EEG-Specific Metadata
Performance Metrics Logging (Limited) ✓ (Yes) (for NLP) ✗ (No) ✓ (Yes, Integrated)
Uncertainty Estimation ✗ (No) ✗ (No) ✗ (No) ✓ (Yes) (Confidence Intervals)
Ethical Considerations ✓ (Encouraged) ✓ (Encouraged) Researcher-Dependent ✓ (EEG Bias Reporting)
Visualization of Results ✗ (No) ✗ (No) ✗ (No) ✓ (Performance Plots)
Reproducibility (Limited, TFX Required) ✗ (No Tracking) ✗ (No Tracking) ✓ (YAML-Based Logging)
Customization (Limited) (Limited to NLP) ✗ (No Format) ✓ (Flexible YAML)
Open Source? ✓ (Yes) ✓ (Yes) ✗ (No) ✓ (Yes)
Table 2: Comparison of Model Documentation Frameworks
@@ 3.1 Software architecture
Figure 2: Workflow of the DREAMS Model Card Generation Process: The diagram illustrates the end-to-end pipeline, starting from data collection and exploration, followed by training and validation phases, and progressing to evaluation metrics and plot generation. Key outputs, including plots and logs, are saved in a designated folder, which are then retrieved by the configuration file (.YAML). Finally, the $generate\_model\_card()$ function utilizes the configuration and saved assets to automatically generate the structured model card.
@@ 3.2.2 Model training and evaluation
At this stage, we meticulously document all the paths to vital information related to both training and evaluation results within the YAML file. This DREAMS comprehensive record encompasses not only paths to graphics but also to other essential data, ensuring that every critical component is easily accessible and well-organized. This approach ensures efficient management and retrieval of information throughout the training and evaluation processes.
@@ 3.2.3 Model card generation
An example of the code snippet using generate_modelcard() function is as follows:
@@
4 Hypothesis and Limitations
The primary hypothesis of this work is that structured documentation of deep learning models through automated model cards can significantly enhance transparency, reproducibility, and ethical compliance in EEG-based deep learning. Specifically, we hypothesize that:
  • A domain-specific model card frameworks (DREAMS) tailored for EEG-based deep learning can improve the reporting of performance metrics, dataset biases, and model uncertainty compared to generic model documentation approaches.
  • Automated documentation of EEG preprocessing steps, dataset details, and evaluation results will reduce inconsistencies and improve reproducibility in deep learning-based EEG research.
  • A standardized model card structure incorporating EEG-specific metadata can facilitate better interpretability and regulatory compliance in healthcare AI applications.
To support this hypothesis, we make some assumptions regarding the use of EEG data and deep learning models:
  • The quality of model documentation is a critical factor affecting the interpretability and trustworthiness of EEG-based deep learning models.
  • The datasets used in this study (FACED and CAUEEG) are representative of real-world EEG classification tasks, allowing generalization of results to similar applications.
  • The reported performance metrics (e.g., accuracy, F1-score, uncertainty estimates) provide meaningful insights into model effectiveness and limitations, which are accurately captured in the generated model cards.
  • Researchers and clinicians will benefit from structured model documentation without significant modifications to existing deep learning workflows.
4.1 Limitations
Despite the advantages of DREAMS, we acknowledge some limitations. The framework’s effectiveness depends on the EEG data quality and preprocessing techniques. Poorly preprocessed or noisy EEG signals may introduce biases that cannot be fully mitigated through documentation alone. Additionally, while DREAMS automates model documentation, the logging and visualization processes introduce computational overhead, which may not be ideal for real-time EEG applications requiring low-latency performance. Another limitation is the framework’s current level of clinical validation; while it is designed for research and development, its direct applicability in clinical environments requires further evaluation with real-world patient data and potential regulatory compliance assessments. Furthermore, although DREAMS is framework-agnostic, its implementation is primarily optimized for Python-based deep learning libraries such as PyTorch and TensorFlow, limiting its immediate usability for researchers working with other AI ecosystems. Lastly, while DREAMS promotes transparency by documenting model biases and performance limitations, it does not eliminate the ethical concerns associated with EEG-based deep learning models. Ensuring fair data collection, privacy protection, and algorithmic fairness remains the responsibility of researchers and developers implementing the framework. By recognizing these challenges, we highlight areas for future improvements and refinements, ensuring DREAMS continues to evolve as a robust and practical model documentation tool for EEG-based AI applications.
5 Illustrative Examples on EEG data
We provide examples of model cards for two models: an EEG-based emotion classifier and an EEG dementia classification model.
5.1 Emotion Classifier
@@
We present a model card for an EEG-based emotion classifier trained on the Finer-grained Affective Computing EEG Dataset (FACED) [14], which is compatible with TorchEEG  [^1] https://torcheeg.readthedocs.io/en/latest/. The dataset consists of EEG recordings from 123 subjects using 32 electrodes positioned according to the international 10–20 system. During the recording sessions, participants viewed 28 different video clips designed to elicit various emotional responses. These video clips spanned nine distinct emotion categories: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral emotion. We applied band-pass filtering, common spatial pattern filtering, and normalization to the EEG data.
Model Card - EEG Emotion Classifier (v 0.1.0)
Overview
This model classifies EEG signals into negative, neutral, and positive classes. It predicts emotions such as anger, disgust, fear, amusement, joy, and more. Additionally, it provides uncertainty estimation of predictions with confidence intervals, hyperparameters, and performance metrics.
Dataset
Dataset: FACED Dataset Number of Classes: 3 Ground Truth: Negative (0), Neutral (1), Positive (2) Training/Validation Split: 80:20 Preprocessing: Band-Pass Filtering, Common Spatial Patterns (CSP), Normalization [Uncaptioned image]
Model Details
Input: 30-channel EEG time-series data Output: Class labels, Confidence intervals Model Type: CNN (Xception-based) Learning Rate: 0.001 Batch Size: 32 Parameters: 0.56M trainable parameters
Performance
The model’s performance is evaluated on an independent test set. Below is a summary of the performance by confusion matrix. [Uncaptioned image]
Limitations
The dataset exhibits class imbalance (male-to-female ratio 3:1), which might introduce bias. Noisy EEG channels and missing data can degrade classification performance.
Uncertainty
Confidence intervals provide insights into model reliability. Below, the performance metrics are reported with 95% confidence intervals. [Uncaptioned image]
Figure 3: An illustrative example of Model Card for EEG Emotion Classification. This model card summarizes the dataset, model details, performance metrics, uncertainity of predictions and limitations for an EEG-based emotion classification system.
Model Card - Abnormal/Normal Classifier (v 1.0.0)
Overview
The model analyzed in this card classifies EEG into abnormal or normal class and returns an uncertainty estimation of the predicted class with a confidence interval. It reports the model’s hyperparameters and performance plots. The reader can gain further insights about the model’s limitation by identifying the input instances where the model is expected to perform well.
Dataset
Dataset: TUH Abnormal Number of Classes: Two Ground Truth: Health status (abnormal or normal) Training/Validation Split: 80:20 Preprocessing: Independent Component Analysis, Band-Pass Filtering, Normalization [Uncaptioned image]
Model Details
Input: 19-channel EEG data in time domain. Output: Class labels abnormal/normal, Confidence intervals Model Type: Brain Signal Vision Transformer(BSVT). Learning Rate: 1e-05 Batch Size: 32 Parameters: 0.75M trainable parameters
Performance
The model’s performance is evaluated on an independent test set. Below is the plot of training and validation accuracy. [Uncaptioned image]
Limitations
Long EEG sequences require large tokenized inputs, leading to excessive memory consumption for the transformer model. Noisy EEG channels and missing data can degrade classification performance.
Uncertainty
Confidence intervals provide insights into model reliability. Below, the performance metrics are reported with 95% confidence intervals. [Uncaptioned image]
Figure 4: An illustrative example of Model Card for EEG Abnormal/Normal Classification. This model card summarizes the dataset, model details, performance metrics, uncertainity of predictions and limitations for the classification of abnormal and normal EEG signals.
@@
We applied a deep learning model called TSception, a multi-scale convolutional neural network designed to classify emotions from EEG data. It incorporates dynamic temporal layers, asymmetric spatial layers and high-level fusion layers, enabling the model to learn discriminative representations across both time and channel dimensions simultaneously. The model described in this example classifies EEG data into three main categories: negative (0), which includes emotions such as anger, disgust, fear, and sadness; neutral (1); and positive (2), which includes emotions like amusement, inspiration, joy, and tenderness.
The model card provides a structured evaluation of model performance, including confusion matrix analysis, from which true positive (TP), false positive (FP), true negative (TN), and false negative (FN) rates are derived to assess classification performance. Additionally, it includes uncertainty estimation, offering insights into the confidence levels of model predictions. Furthermore, the model card provides comprehensive documentation, outlining key aspects of the dataset characteristics, preprocessing steps applied to EEG signals, and an overview of the deep learning model for emotion classification.
Figure 3 illustrates the prototype of the model card, highlighting its structured approach to model transparency and interpretability. Notably, the model card is fully customizable and can be extended to incorporate additional domain-specific details, depending on the specific application requirements.
5.2 Abnormal Classifier
We provide a second illustrative model card using the publicly available abnormal EEG dataset from the Temple University Hospital (TUH) v2.0.0 [15]. The dataset includes 2,993 recordings from 2,329 different subjects. The recordings are labelled as normal (1,385 recordings) or abnormal (998 recordings). The training set contains 2,717 recordings and the evaluation set holds 276 recordings. We applied band pass filtering, Independent Component Analysis (ICA), Normalization as the preprocessing steps on the input dataset.
A transformer-based model, BSVT [16], was employed for the classification task, with its performance and prediction uncertainty systematically documented in the model card. Figure 4 presents an example of the model card, showcasing how variations in model architecture and performance evaluation can be structured within a standardized format. Notably, if hyperparameters are modified or an alternative model is selected, the model card can accommodate these updates by specifying the version number and associated changes. Furthermore, the model card is highly customizable, allowing for the inclusion of additional sections as needed. For instance, a recommendation section could be added to provide additional information on model deployment, usage considerations, or the necessity for further validation and testing.
5.3 Pipeline for Model Card Generation
This section details the step-by-step process for training a deep learning model using pytorch framework, evaluating its performance, and ultimately generating a structured model card. The pipeline includes data preprocessing, model training, validation, performance analysis, and automated documentation using a model card.
Data Preprocessing and Preparation The EEG data undergoes online transformation into tensor format and a 2D representation for compatibility with deep learning models. The labels are adjusted accordingly, and the dataset is split into training and validation sets. To efficiently manage data batching and shuffling, we create DataLoaders for both training and validation. The implementation is as follows:
@@ 5.3 Pipeline for Model Card Generation
3dataset = Dataset(root_path= data_folder,
@@
Model Setup and Configuration At this stage of the pipeline, the model is initialized and configured with the necessary parameters. The model is instantiated using the Model function, specifying key hyperparameters for instance, in our case the number of classes, electrodes, sampling rate, and architectural components. The following code snippet demonstrates the initialization:
1
2model = Model(num_classes=9,num_electrodes=30,
3 sampling_rate=250,num_T=15,
4 num_S=15, hid_channels=32, dropout=0.5)
Training and Validation Process We proceed to configure the training and validation processes utilizing PyTorch. This involves setting up the necessary parameters and components for training the model, such as defining the loss functions, optimizers, and evaluation metrics. We establish the procedures for validating the model’s performance, including managing the data flow, monitoring metrics, and adjusting the training process as needed to ensure effective learning and evaluation. A standard implement for the training and validation configuration is:
@@
Performance Evaluation and Metrics Visualization To assess the model’s performance, we generate a confusion matrix and compute key evaluation metrics, such as accuracy, precision, recall, and F1-score. Additionally, uncertainty estimation is applied to quantify the confidence of predictions.
@@
3plot_confidence_intervals(precision=results[’precision’], recall=results[’recall’] ...
Model card Generation Lastly, we call the generate_model_card() function, providing it with the path to the YAML configuration file  [^2] https://github.com/LucidJun/DREAMS/tree/main/template, the output path for saving the model card, and the version number of the model card as shown below:
@@
Figure 3 and 4 provide the illustrative example of the generated model cards following the above describe pipeline. This model card encapsulates the key information and metrics resulting from the model’s training and evaluation processes, offering a comprehensive summary of the model’s performance, data characteristics, and other relevant details.
6 Impacts